OpenStudy (anonymous):

What is the range of the function f(x) = 12 − 3x for the domain {-4, -2, 0, 2, 4}?

OpenStudy (anonymous):

@Haseeb96

OpenStudy (haseeb96):

just put the one by one , each element of set domain into the given function and get the range for each element.

OpenStudy (haseeb96):

like if the function is this f(x) = 2x +1 and domain is 2 now find the range i will just put the domain at the place of x in the given function f(2)= 2(2)+1 = 4 +1 = 5 so 5 is your range

OpenStudy (haseeb96):

now solve it like this and get the range

OpenStudy (haseeb96):

@Alexandra123456 now solve it yourself and show me your answers

OpenStudy (anonymous):

i dont get it i get decmals and it is not the correct answer @Haseeb96

OpenStudy (haseeb96):

okay let me solve it first put -4 in the function at the place of x f(-4)= 12 -3(-4) =12 +12 =24 24 is the range of -4 now solve it yourself

OpenStudy (anonymous):

okay so the answer would be [0,6,12,18,24] thats what i get

OpenStudy (anonymous):

@Haseeb96

OpenStudy (haseeb96):

absolutely correct. Good job smart girl :)

OpenStudy (haseeb96):

@Alexandra123456

OpenStudy (anonymous):

Thanks! @Haseeb96

OpenStudy (haseeb96):

Most welcome my new user :)